Scores on a standardized test are normally distributed with a mean of 246 and a standard deviation of 22. Students who score at least 2.4 standard deviations above the mean receive a certificate. What is the minimum score required to receive the certificate?
Answer by stanbon(75887)

Scores on a standardized test are normally distributed with a mean of 246 and a standard deviation of 22. Students who score at least 2.4 standard deviations above the mean receive a certificate. What is the minimum score required to receive the certificate?
-------------------------
Use the form x = z*s+u
---
x = 2.4*22+246 = 298.8
===================
Cheers,
Stan H.
===================